a guy has coffee worth $60 a pound that he wishes to mix with 50 pounds of coffee worth $90 a pound to get a mixture that can b sold for $80 a pound. How many pounds of the $60 coffee should she use
Answer by stanbon(75887)

a guy has coffee worth $60 a pound that he wishes to mix with 50 pounds of coffee worth $90 a pound to get a mixture that can b sold for $80 a pound. How many pounds of the $60 coffee should she use
------
Equation:
value + value = value
60x + 90*50 = 80(x+50)
60x 4500 = 80x + 4000
-20x = -500
x = 25 lbs (amt. of $60 coffee needed)
========================================
Cheers,
Stan H.
===========